Ian Pearson reports that Tuesday was a lovely fine day at Grosmont with Nigel Bill, Ian McCall, Neil Smedley, Nigel Hall, Jon Bradley, Gordon Wells and Ian himself.
A successful day with the following achieved:
- Lefthand piston packing fitted and secured with slide bar oil pot attached.
- Lefthand little end pin fitted and secured with castellated nut and cotter fitted.
- Lefthand and righthand big end felt pads fitted.
- Lefthand and right hand big end securing nuts fitted and tightened, then taper pin fitted.
- Jon Bradley steamed a wood strip for fitting round the edge of the cab roof (see photo).
- Righthand crosshead cotter fitted but securing cotter to make and fit on Wednesday.
- Righthand trailing coupling rod end cap fitted and taper pinned
- Lefthand top and bottom eccentric rods fitted. Some securing still to do with tightening of nuts on to the eccentric straps and taper pins to secure and open out at the link end
- The large lefthand crosshead cotter was hammered down but jammed part way in, so this will require removal next week.
- A hose was linked up to the main water supply and fed into the tender and filling commenced
- The tender was measured up on the righthand side ready for the BR transfer to be fitted.
- Loco number transfers 63395 put on the cab sides (see photo) and then both cab sides varnished.
Paul and Angie arrived at about 17:30 with a washout pump. They exchanged that for seat cutters.
Wednesday was another fine day warm day at Grosmont until later in the afternoon when the heavens opened. Hopefully that will have reduced the potential fire risk on the moors at Summit. Only Nigel Hall and Steve Hyman were present in Deviation Shed although Ian called in to see how things were going, and Peter Whitaker came and put the lion portion of the tender transfer on, leaving the two British Railway pieces for Nigel to complete (see photo).
Nigel Hall did a lot of tidying up around Deviation Shed, sweeping the walkways and moving items to clear the walkway, especially between Nos 6 and 7 Roads as there will be a Vera Lynn working part on Saturday with Ian Foot. Steve completed a number of jobs around the Q6's motion:
- Made a small securing cotter and fitted to the righthand large cotter.
- Small cotter made and fitted to the intermediate side rod gradient pin large cotter.
- Lefthand little end oil pot fitted.
- Lefthand piston oil pot fitted.
Nigel continued painting control handles etc in the afternoon after Steve had left, as it was unlikely that anybody was going to access the cab or tender before the paint had dried.
The tender was topped up with the water level about 18” from the top of the tank. All filling hoses have been returned to the MPD hose lockers inside No 5 Road. Filling was required to prevent the tender from derailing when shunting, as the Q6 is due to be moved into the running shed tomorrow (Friday). Both sides of the tender now need a coat of varnish to seal the transfers and Nigel Hall was hoping to do that before the Q6 is moved. The machined steam pipes were collected from Oak Engineering at Hebburn by Nigel Bill yesterday afternoon and delivered to Grosmont this morning. They will be fitted by the MPD staff. The gauge glasses, Klinger packings and grate need fitting, but we are now getting very close to an internal steam test prior to that for the boiler inspector, and a return to traffic.